When two idenical batteries of int. resistance 1ohmeach are connected in series across a resistor R is J1. When the same batteries are connected in parreel across R the rate is J2. If J1=2.25J2 the value of R in ohms is-

Grade:12

1 Answers

Vikas TU
7 years ago
in series,
J1 = 4E^2/(R+2)
and
in parallel,
net emf = (E/1 + E/1)/(1/1 + 1/1) = > 2E/2 = E
J2 = E^2/R
 
Now it is given J1 = 2.25J2
put  both J1 and J2 values and get the value of R.